an organic chemist isolates a colorless liquid from a petroleum sample she places the liquid in preweighed flask and put the flask in boiling water which vaporizes the liquid and fill the flask with gass she closes the flask and reweighs it she obtained the following data
volume of flask =213ml
T=100.0 degree celcius , p= 754
mass of flask+gass=78.834g
mass of flask = 77.83 g
calculate molar mass of liquid
1
Expert's answer
2020-12-14T14:41:57-0500

V1 = 213mL

T1 = 100°C = 373K

P1 = 754mmHg

at STP

V1 = ?

T2 = 0°C = 273K

P2 = 760mmHg


Using the general gas law;


"\dfrac{P1V1}{T1} = \dfrac{P2V2}{T2}"


"V2 = \dfrac{P1V1T2}{P2T1}"


"V2 = \dfrac{754×213×373}{760×373} = 211.3mL"



mass of flask + gas = 78.834g

mass of flask = 77.83g

therefore, mass of gas = 1.004g


1.004g = 211.3ml

xg = 22400ml (molar volume of gas)


x = "\dfrac{22400×1.004}{211.3}" = 106.43g


the gas with the neatest molar mass of 106.43g is Palladium.


Therefore, the liquid heated is Palladium with a molar mass of 106.42g and an atomic number of 46.